OpenStudy (anonymous):

Please Help! A pizza restaurant, Pizz-A, plans on selling 115 pizzas daily for delivery. For every 15 customers who eat in the restaurant, an additional 4 pizzas will be sold. Part A: Write an equation to represent how many pizzas a day Pizz-A plans on selling in terms of the number of eat-in customers. Identify the meaning of all variables used. Part B in comments

OpenStudy (anonymous):

A pizza restaurant, Pizz-A, plans on selling 115 pizzas daily for delivery. For every 15 customers who eat in the restaurant, an additional 4 pizzas will be sold. Part A: Write an equation to represent how many pizzas a day Pizz-A plans on selling in terms of the number of eat-in customers. Identify the meaning of all variables used. Part B: What would an increase in the y-intercept represent? Part C: Create a second equation for an eat-in only restaurant, in the same chain, with a lower ratio of pizzas sold to eat-in customers. Does this equation have the same intercept and slope as the equation from part A? Explain your reasoning.

OpenStudy (anonymous):

*15x = 4y

OpenStudy (anonymous):

A: y=115+4(n/15)

OpenStudy (anonymous):

Do you understand the equation I gave you?

Directrix (directrix):

I am not saying that this is correct. The following comes from Brainly.com http://brainly.com/question/133689 Work by @Anthonybayley1 A: y=115+4(n/15) where n is the number of customers. y=115+4n/15

OpenStudy (anonymous):

Part A: Write an equation to represent how many pizzas a day Pizz-A plans on selling in terms of the number of eat-in customers. Identify the meaning of all variables used. Part A: \[ y = \frac{4}{15}x + 115\] y = amount of pizza sold. x = number of customers. Part B: What would an increase in the y-intercept represent? The increase in the y-intercept would represent that Pizz-A would be planning to sell more pizza by delivery. Part C: Create a second equation for an eat-in only restaurant, in the same chain, with a lower ratio of pizzas sold to eat-in customers. Does this equation have the same intercept and slope as the equation from part A? Explain your reasoning. Create a second equation for an eat-in only restaurant, in the same chain, with a lower ratio of pizzas sold to eat-in customers. Part C Equation: \[ y = \frac{1}{4}x \] Part C: Does this equation have the same intercept and slope as the equation from part A? Explain your reasoning. No it does not have the same intercept because this is an eat-in only restaurant and no sales are expected from delivery and no sales are predicted so the y-intercept will be 0 for the new equation. It also does not have the same slope because the ratio of pizza sold to eat-in customers is lower. The new ratio is 1 pizza is sold ever 4 customers. Work By @MDoodler

OpenStudy (anonymous):

Thanks @MDoodler